OpenStudy (anonymous):

HELP:( Would someone mind helping the function f(x) varies inversely with x and f(x)=8 and when x=3. what s f(x) when x=2

OpenStudy (victoriasushchik):

First, let y=k/x. Now, we know that y=12 when x=3, so we can solve for k: 12=k/3, or k=12*3=36. Now we have y=36/x. So y=36/20 when x=20. Thus, y= 9/5. Hope this helps!

OpenStudy (anonymous):

whats 9/5?

OpenStudy (anonymous):

theres no answer for that on my study sheet

OpenStudy (anonymous):

x=2 not 20

OpenStudy (victoriasushchik):

THATS THE STEPS HOW TO DO THEM

OpenStudy (anonymous):

oh well i am confused. not very good at mathh

OpenStudy (anonymous):

so would it be 12?